OpenStudy (anonymous):

How do you solve? N - .0667N = 9880.48

OpenStudy (ajprincess):

First take N as the common factor in RHS. You wil get N(1-0.667)=9880.48 0.333N=9880.48 Divide by 0.333 both sides. N=9880.48/0.333 =29671.111